OpenStudy (anonymous):

Please Help!!! Two cars are driving toward each other from opposite directions. If one is driving 55 mph, the other is driving 60 mph and the cars are 500 miles apart, how long will it be before they pass each other? Round to the nearest tenth of an hour. A. 4.3 hours B. 4.5 hours C. 5.1 hours D. 4.8 hours

OpenStudy (anonymous):

60*t + 55*t = 500 => t = 500/115 => t = 4.3 hours.

OpenStudy (dumbcow):

d=r*t car1: d1 = 55t car2: d2 = 60t d1 +d2 = 500 --> 55t+60t = 500 --> 115t = 500 --> t = 500/115